OpenStudy (anonymous):

identify the open intervals on which the function is increasing or decreasing-- f(x)=(x+3)^3

OpenStudy (tkhunny):

Do we get to use the derivative?

OpenStudy (anonymous):

yup!

OpenStudy (tkhunny):

All-righty, then. Show us f'(x).

OpenStudy (anonymous):

3(x+1)^2

OpenStudy (shamil98):

the derivative is 3(x+3)^2

OpenStudy (tkhunny):

Perfect. If you think of the right thing, the answer to the following question is really, REALLY easy. :-) Where is f'(x) negative, given that \(f'(x) = 2(x+3)^{2}\)?

OpenStudy (anonymous):

wait did i get the derivative wrong

OpenStudy (tkhunny):

No, you got it. I just had a typo-spasm. The answer to my question is the same. When is f'(x) negative? Don't look at the graph. Just think on the structure. A Real Number squared. When is that negative?

OpenStudy (shamil98):

I got a question, why is 3(x+3)^2 not the derivative? o.o

OpenStudy (anonymous):

Isnt it using the chain rule...maybe haha

OpenStudy (tkhunny):

?? Why as that? You have it. Don't let my typo confuse you. \(f(x) = (x+3)^{3}\) \(f'(x) = 3(x+3)^{2}\) Okay, now answer... When is f'(x) negative?

OpenStudy (anonymous):

so actually my question was f(x)=(x+1)^3 but thats okay haha but im not sure it can be negative

OpenStudy (anonymous):

idk

OpenStudy (shamil98):

oh then the derivative is 3(x+1)^2

OpenStudy (anonymous):

hahaha ya

OpenStudy (tkhunny):

Come on. You can see it. If you start with a Real Number, and Square it, will you EVER get a negative number? The derivative is NOT \(3(x+1)^{2}\). The derivative is \(f'(x) = 3(x+1)^{2}\). Don't be afraid to write whole, complete expressions.

OpenStudy (anonymous):

no

OpenStudy (shamil98):

sorry, f'(x) = 3(x+1)^2 continue with your explanation :P

OpenStudy (tkhunny):

Can f'(x) EVER be zero (0)?

OpenStudy (anonymous):

if x was -1 right?

OpenStudy (anonymous):

haha im stupid at calc. sorry :(

OpenStudy (anonymous):

tkhunny where did you go????